OpenStudy (anonymous):

find the area for a triangle with the bases b and the height h? b=7 cm h=12 cm

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

\[\rm Area = \dfrac{1}{2 } \times base \times height \]

Parth (parthkohli):

Can you continue?

OpenStudy (anonymous):

is it 42

Parth (parthkohli):

lol, of course it is! ^_^ Don't you know that the answer to everything, the life and the universe is 42? :-D

OpenStudy (anonymous):

ok thanks